OpenStudy (anonymous):

Determine the x- and y- intercepts for the graph defined by the given equation. y = 7x + 3

OpenStudy (sidsiddhartha):

try to write it in \[\frac{ x }{ a }+\frac{ y }{ b }=1 \] form then a=x intercept and b=y intercept

OpenStudy (anonymous):

?? What im confused

OpenStudy (sidsiddhartha):

ok u can write it as \[7x-y=-3\] right?

OpenStudy (sidsiddhartha):

now divide both sides by (-3) \[\frac{ 7x-y }{ -3 }=-3/-3\\ \frac{ -7 }{ 3 }x+\frac{ y }{ 3 }=1\] getting this?

OpenStudy (sidsiddhartha):

so finally it can be written as \[\large\frac{ x }{ \frac{ -3 }{ 7 } }+\frac{ y }{ 3 }=1\] so x intercept:-3/7 y intercept:3

OpenStudy (sidsiddhartha):

ok ?

OpenStudy (anonymous):

Determine the x- and y- intercepts for the graph defined by the given equation. x = -9 a. x-intercept is ( -9, 0) No y-intercept c. No x-intercept y-intercept is (0, -9) b. No x-intercept y-intercept is ( -9, 0) d. x-intercept is (0, -9) No y-intercept